OpenStudy (anonymous):

Factor, check by multiplying factors ac-3ad-2bc+6bd

OpenStudy (anonymous):

a(c-3d) -2b(c-3d) = (a-2b)(c-3d)

OpenStudy (anonymous):

\(\Large ac-3ad-2bc+6bd=(ac-2bc)-(3ad-6bd)\\\Large =c(a-2b)-3d(a-2b)=(a-2d)(c-3d)\)